The table below shows the marks scored by

229

students in a statistics class test.


Marks

10  20 20  30 30  40 40  50 50  60 60  70 70 80

Frequency

12 30 x 65 y 25 17

 Given that the median mark is

46

, determine the:

(i) Values of

x

and

y

; [5 Marks]

(ii) Mode; [2 Marks]

(iii) Mean absolute deviation; [4 Marks]

(iv) Quartile deviation. [4 Marks]


1
Expert's answer
2021-11-30T16:48:34-0500

i)

for grouped data:

median = "L+\frac{w(n/2-B)}{G}"

where:

  • L is the lower class boundary of the group containing the median
  • n is the total number of values
  • B is the cumulative frequency of the groups before the median group
  • G is the frequency of the median group
  • w is the group width


median group: 40-50

cumulative frequency of median class: "B=12+30+x"

then:

median = "46=40+\frac{10(229/2-(12+30+x))}{62}"


"x=229/2-6\cdot62/10-12-30=35"

"y=n-12-30-35-65-25-17=229-12-30-35-65-25-17=45"


ii)

Mode = "L + \frac{f_m − f_{m-1}}{(f_m − f_{m-1}) + (f_m − f_{m+1})} w"

where:

  • L is the lower class boundary of the modal group
  • fm-1 is the frequency of the group before the modal group
  • fm is the frequency of the modal group
  • fm+1 is the frequency of the group after the modal group
  • w is the group width


modal group: 40-50


Mode = "40 + \frac{10(65 − 35)}{65-35 +65-45} =46"


iii)

Mean absolute deviation:

"dm=\frac{\sum |x_i-\overline{x}|f_i}{n}"

where fi is frequency,

xi is midpoint of class,

mean:

"\overline{x}=\frac{\sum f_ix_i}{n}=\frac{15\cdot12+25\cdot30+35\cdot35+45\cdot65+55\cdot45+65\cdot25+75\cdot17}{229}=45.7"


"dm=\frac{30.7\cdot12+20.7\cdot30+10.7\cdot35+0.7\cdot65+9.3\cdot45+19.3\cdot25+29.3\cdot17}{229}=12.26"


iv)

 Quartile deviation:

"QD=(Q_3-Q_1)\/2"


Qi class = (in/4)th value of the observation


"Q_i=L+\frac{in/4-cf}{f}⋅w"

where cf is cumulative frequency


Q1 class = (229/4)=(57.25)th value of the observation

class: 30-40


"Q_1=30+\frac{57.25-77}{35}⋅10=24.36"


Q3 class = ("229\cdot3/4" )=(171.75)th value of the observation

class: 50-60


"Q_1=50+\frac{171.75-187}{45}⋅10=46.61"


"QD=(46.61-24.36)\/2=11.13"